NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 18:55:36 on June 13, 2009
Entry number 281092
Followups:
keyword=Experiment Summary for June 11 owl - June 12 swing
Summary for June 11 owl - June 12 swing
Completed 2nd pass production running and switched to
3rd pass.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/01/11 owl | 6.14 |
0.42 | 1.43 |
2009/01/11 day | 0.76 |
0.18* | 7.06* |
2009/01/11 swing | 3.27 |
2.47* | 2.26* |
2009/01/12 owl | 6.50 |
0.48 | 1.04 |
2009/01/12 day | 0.00 |
1.08 | 6.92 |
2009/01/12 swing | 3.96 |
2.15 | 1.90 |
total | 20.6 |
6.8 | 20.6 |
* These numbers are corrected compared to the shift leader submitted
numbers.
Main events/issues:
June 11, 2009
- BHWP: IN, Target spin: Longitudinal
- Continue 3He(e,e'd) production for left HRS.
- 00:05 Set right HRS P0 = 0.7 GeV/c for PVDIS test plan.
- 02:27-02:46 Beam down due to IOCHA1.
- 03:00 Right dipole finally set.
- 09:10-17:45 Contolled Access:
- Neutron detector moved to 54o and 8 m from target.
- Installed right and left HRS sieve-slits.
- Removed Deuterium bottle and connected hydrogen gas bottle.
- Compton work.
- Right momentum set to 2.175 GeV/c.
- 10:00-17:00 Beam studies and shed load power test.
- 18:00-19:32 HRS optics studies.
- 19:35-20:10 Contolled access:
- Remove HRS sieve-slits.
- Insert BigBite sieve-slit.
- 20:10-22:30 BigBite optics studies.
- 22:32-23:00 Controlled access to remove BigBite sieve-slit.
- 23:00-24:00 Reference cell data: empty, N2 and
3He.
- Average target polarization for June 11th: 62.1%.
June 12, 2009
- BHWP: IN, Target spin: Longitudinal
- 00:30-07:35 Continue 3He(e,e'd) production for left HRS.
- Begin production for right HRS in dip region: P0 = 2.025
GeV/c.
- 07:35-08:46 QE measurement of cathode and spot move.
- 08:00 Change right and left P0 = 3.0855 GeV/c.
- 08:33-09:55 Controlled access to move spectrometers:
- Right HRS moved to 17o.
- Left HRS bogie controls unresponsive. Decided to leave at
12.5o.
- Left HRS P0 set to 3.35 GeV/c.
- 09:30-16:00 Pass change from 2nd to 3rd.
- 10:45-12:05 Controlled access to rest HV crates.
- 16:00-18:25 Beam tuning and profile check out.
- 18:25 Begin 3rd pass production at 10 uA.
- Right HRS: QE 3He(e,e'n) for
GEn measurement.
- Left HRS: QE 3He(e,e'), 3He(e,e'p) and
3He(e,e'd) measurements.
- 20:28-21:08 Controlled access for Compton work.
- 21:26 Resume production on polarized 3He.
- 23:50 Right HRS dipole appears to have quenched taking right Q1,
Q3 and
left Q1 and the dipole out with it.
- Waiting for liquid levels to restore to normal values.
- Average target polarization for June 12th: 64.7%.